# Structure and Independence in Hyperbolic Uniform Disk Graphs

### v1

@@ 1 Introduction
Our algorithm yields the first component of the running time ($n^{\mathcal{O}\big{(}1+\frac{1}{r}\log k\big{)}}$), but for very small $r$ we can switch to the general disk graph algorithms with running times $2^{\mathcal{O}(\sqrt{n})}$ by De Berg et al. [BergBKMZ20] and $n^{\mathcal{O}(\sqrt{k})}$ by Marx and Pilipczuk [MarxP22]. Our algorithm is best possible under ETH for $r=\Theta(1)$ as [Kisfaludi-Bak20] showed a lower bound of $n^{\Omega(\log n)}$ (for large $k$). Notice moreover that the almost-Euclidean case of $r=1/\sqrt{n}$ gives a running time of $n^{\mathcal{O}(\sqrt{n}\log k)}$, which almost recovers the Euclidean running time of $2^{\mathcal{O}(\sqrt{n})}$. Recall that the Euclidean running time is also ETH-tight for Euclidean unit disks [BergBKMZ20]. While this Euclidean lower bound cannot be directly applied to $\mathrm{HUDG}(1/\sqrt{n})$, it can be adapted to the this setting, so the running time lower bound $2^{\Omega(\sqrt{n})}$ holds also in the hyperbolic plane (assuming ETH). See [Kisfaludi-Bak21] for a similar adaptation of a Euclidean lower bound to the setting of $r=1/\sqrt{n}$.
Finally, but most surprisingly, we get a polynomial running time for Independent Set in the firmly hyperbolic setting of $r\in\Omega(\log n)$; in fact, our algorithm is polynomial already for $r\in\Omega(\log k)$. In particular, this provides a polynomial exact algorithm for hyperbolic random graphs. It thereby directly generalizes the algorithm of [BlasiusFFK23] which gives a polynomial running time in hyperbolic random graphs with high probability: our algorithm has no assumptions on the input graph distribution and provides a polynomial worst-case running time.
The underlying idea of our exact algorithm can be regarded as a dynamic programming algorithm along the (unknown) tree decomposition of the Voronoi diagram of the solution disks, inspired by Marx and Pilipczuk [MarxP22]. More precisely, both [MarxP22] and the present paper use so-called sphere cut decompositions, a variant of branch decompositions for plane graphs [Dorn_Penninkx_Bodlaender_Fomin_2010] to guide the algorithm. We introduce well-spaced and valid sphere cut decompositions, and prove that they are in one-to-one correspondence with independent sets of $G$.
Finally, we consider approximation algorithms for Maximum Independent Set when the graph in question has low ply, that is, each point of $\mathbb{H}^{2}$ is covered by at most $\ell$ disks. In Section 6 we show the following.
Theorem 4.
@@
An important component of the algorithm’s correctness proof is bounding the so-called degeneracy of hyperbolic uniform disk graphs in terms of their clique size and in terms of their ply. This generalizes the Euclidean results of [MaratheBHRR95]. When compared to the $n^{\mathcal{O}(1/\varepsilon)}$ algorithm for disk graphs by Chan [Chan03] or the $2^{\mathcal{O}(1/\varepsilon)}n$ algorithm in planar graphs by Baker [Baker94], both of which are conditionally optimal [Marx07a], our algorithm has a surprising quasi-polynomial dependence on $1/\varepsilon$ rather than exponential. The dependence on $r$ exhibits the classic Euclidean-to-hyperbolic scaling seen in our earlier structural results. For $\varepsilon=1/n$, we are guaranteed to get the optimal solution, and setting $\ell=n$ covers the general case. The resulting running time for $1/\varepsilon=\ell=n$ is $n^{\mathcal{O}\left(1+\frac{1}{r}\log n\right)}$, which matches our exact algorithm. Consequently, our approximation scheme is a generalization of our exact algorithm.
@@ 2 Preliminaries
An independent set in a graph is a set of pairwise non-adjacent vertices. In the Independent Set problem, we are given a graph $G$ [^6] In this article, the graphs are always intersection graphs given through their geometric representation. and a number $k$, and the goal is to decide if there is an independent set of size $k$. In Maximum Independent Set, the input is a graph $G$ and a number $\varepsilon>0$, and the goal is to output an independent set of size at least $(1-\varepsilon)k^{*}$, where $k^{*}$ is the size of the maximum independent set.
Let $G=(V,E)$ be a graph. A vertex set $S\subseteq V$ is a separator if $V$ can be partitioned into non-empty subsets $S$, $V_{1}$, and $V_{2}$ such that there is no edge between $V_{1}$ and $V_{2}$, i.e., for every $v_{1}\in V_{1}$ and $v_{2}\in V_{2}$ it holds that $\{v_{1},v_{2}\}\notin E$. The separator is $\beta$-balanced if $|V_{1}|,|V_{2}|\leq\beta n$ for some $\beta<1$.
@@
We will need the notion of weighted treewidth [EijkhofBK07]. Here each vertex has a weight, and the weighted width of a tree decomposition is the maximum over the bags of the sum of the weights of the vertices in the bag (note: without the minus $1$). The weighted treewidth of a graph is the minimum weighted width over its tree decompositions. Now let $\mathcal{P}$ be a clique partition of a given graph $G$. We apply the concept of weighted treewidth to $G_{\mathcal{P}}$, where we assign each vertex $C$ of $G_{\mathcal{P}}$ the weight $\log(|C|+1)$, and refer to this weighting whenever we talk about the weighted treewidth of a contraction $G_{\mathcal{P}}$. For any given $\mathcal{P}$, the weighted treewidth of $G_{\mathcal{P}}$ with the above weighting is referred to as the $\mathcal{P}$-flattened treewidth of $G$.
@@
A polygon is a cycle together with a planar drawing in which all edges are drawn as line segments.
@@
Figure 3: Part (a) shows a visualization of the Poincaré disk including (parallel) lines, a circle with its center, a triangle and generalized polygons. Part (b) shows a visualization of the Beltrami-Klein model including (parallel) lines, a triangle, and a generalized polygon. Part (c) shows a set of sites $S\subseteq\mathbb{H}^{2}$ in the Poincaré disk model, $\mathcal{D}(S)$ (black), and $\mathcal{V}(S)$ (blue); Voronoi vertices are filled squares, ideal Voronoi vertices are empty squares.
@@
We write $\mathbb{H}^{2}$ to refer to the hyperbolic plane and $\mathbb{R}^{2}$ to refer to the Euclidean plane. The Poincaré disk is a model of $\mathbb{H}^{2}$ that maps the whole hyperbolic plane into the interior of a Euclidean unit disk; see 3(a). We refer to the center of the Poincaré disk as origin. Straight lines in $\mathbb{H}^{2}$ are represented as circular arcs perpendicular to the boundary of the Poincaré disk or as chords through the origin. Hyperbolic circles are represented as Euclidean circles. The center of the hyperbolic circle is, however, farther from the origin than the Euclidean center of its representation. The Poincaré disk is conformal, i.e., angle preserving.
@@
Points on the boundary of the Poincaré disk are called ideal points. These are not part of the hyperbolic plane, but form a boundary of infinitely far points. An ideal arc between two ideal points $q_{1}$ and $q_{2}$ is the set of ideal points between $q_{1}$ and $q_{2}$ when moving clockwise on the boundary of the Poincaré disk. A generalized polygon is a cycle together with a planar drawing that maps each vertex to a point in $\mathbb{H}^{2}$ or to an ideal point. Each edge is either a line segments between points, a ray from a point to an ideal point, a line between two ideal points, or an ideal arc between two ideal points. Note that the vertices do not completely determine the generalized polygon, as two ideal points can be connected via a line or an ideal arc. Moreover, a 2-cycle can yield a generalized polygon by mapping both vertices to ideal points and one edge to the line and the other to the ideal arc between the two. Figure 3(a) shows two generalized polygons with their interior shaded in blue.
@@
For a straight line $\ell$ and a point $p\notin\ell$, there are infinitely many parallel lines through $p$, i.e., lines through $p$ that do not intersect $\ell$. Two of these parallel lines are special in the sense that they are the closest to not being parallel. They each share an ideal endpoint with $\ell$ and are called limiting parallels. Let $q\in\ell$ be such that $pq$ is perpendicular to $\ell$. Then the angle between $pq$ and the two limiting parallels is the same on both sides and only depends on the length $x=|pq|$. It is called the angle of parallelism and denoted by $\Pi(x)$. It holds that $\sin(\Pi(x))=1/\cosh(x)$ [Euclid_Non_Euclid_Geomet-Green93, page 402].
@@
Let $a$ be the sum of interior angles of a hyperbolic triangle. Then $a<\pi$ and the area of the triangle is $\pi-a$. Note that this implies that the area of a triangle is upper bounded by $\pi$. More generally, the area of a $k$-gon is bounded by $(k-2)\pi$. The area of a disk with hyperbolic radius $r$ is $4\pi\sinh^{2}(r/2)$. For $r\to\infty$ this area is in $\Theta(e^{r})$ and for $r\to 0$ it is in $\Theta(r^{2})$.
@@
The Beltrami–Klein model also uses the interior of the Euclidean unit disk as ground space. Hyperbolic straight lines are represented as chords of the unit disk, see Figure 3(b). This model is not conformal, but enables an easy translation of Euclidean line arrangements into the hyperbolic plane.
@@
Let $S$ be a set of sites, i.e., a set of designated points in $\mathbb{H}^{2}$. The Voronoi cell of a site $s\in S$ is the set of points that are closer to $s$ than to any other site. When considered in the Poincaré disk, the boundary of each Voronoi cell is a generalized polygon [^7] We note that Voronoi cells containing ideal points in their boundary are actually unbounded in the hyperbolic plane.. The non-ideal segments are part of the perpendicular bisector of two sites and are called Voronoi edges; they are illustrated in blue in Figure 3(c). The points in which three or more Voronoi edges meet are called Voronoi vertices. The ideal points in which unbounded Voronoi edges end are called ideal Voronoi vertices. We define the Voronoi diagram $\mathcal{V}(S)$ of $S$ as the following plane graph. Its vertex set is the set of all (ideal) Voronoi vertices. The edge set of $\mathcal{V}(S)$ is comprised of the Voronoi edges and the set of ideal arcs connecting consecutive ideal vertices (red in Figure 3(c)).
The weak dual of the Voronoi diagram is called Delaunay complex and denoted by $\mathcal{D}(S)$; it is illustrated in black in Figure 3(c). The edges of $\mathcal{D}(S)$ are exactly the edges dual to the Voronoi edges, i.e., the edges of $\mathcal{V}(S)$ that are not ideal arcs. Thus, sites $s_{1}\in S$ and $s_{2}\in S$ are connected in the $\mathcal{D}(S)$ if and only if their Voronoi cells share a boundary. In contrast to the Euclidean case, the outer face of $\mathcal{D}(S)$ is not the convex hull of $S$. In fact, its boundary is not necessarily a simple cycle; see Figure 3(c). A site is an outer vertex of $\mathcal{D}(S)$ if and only if the corresponding Voronoi cell is unbounded.

### v2

@@ 1 Introduction
Our algorithm yields the first component of the running time ($n^{\mathcal{O}\big{(}1+\frac{1}{r}\log k\big{)}}$), but for very small $r$ we can switch to the general disk graph algorithms with running times $2^{\mathcal{O}(\sqrt{n})}$ by the authors in [15] and $n^{\mathcal{O}(\sqrt{k})}$ by Marx and Pilipczuk [37]. Our algorithm is best possible under ETH for $r=\Theta(1)$ as [27] showed a lower bound of $n^{\Omega(\log n)}$ (for large $k$). Notice moreover that the almost-Euclidean case of $r=1/\sqrt{n}$ gives a running time of $n^{\mathcal{O}(\sqrt{n}\log k)}$, which almost recovers the Euclidean running time of $2^{\mathcal{O}(\sqrt{n})}$. Recall that the Euclidean running time is also ETH-tight for Euclidean unit disks [15]. While this Euclidean lower bound cannot be directly applied to $\mathrm{HUDG}(1/\sqrt{n})$, it can be adapted to the this setting, so the running time lower bound $2^{\Omega(\sqrt{n})}$ holds also in the hyperbolic plane (assuming ETH). See [28] for a similar adaptation of a Euclidean lower bound to the setting of $r=1/\sqrt{n}$.
Finally, but most surprisingly, we get a polynomial running time for Independent Set in the firmly hyperbolic setting of $r\in\Omega(\log n)$; in fact, our algorithm is polynomial already for $r\in\Omega(\log k)$. In particular, this provides a polynomial exact algorithm for hyperbolic random graphs. It thereby directly generalizes the algorithm of [4] which gives a polynomial running time in hyperbolic random graphs with high probability: our algorithm has no assumptions on the input distribution and provides a polynomial worst-case running time.
The underlying idea of our exact algorithm can be regarded as a dynamic programming algorithm along the (unknown) tree decomposition of the Voronoi diagram of the solution disks, inspired by Marx and Pilipczuk [37]. More precisely, both [37] and the present paper use so-called sphere cut decompositions, a variant of branch decompositions for plane graphs [17] to guide the algorithm. However, due to our setting, a simple adaptation of the divide-and-conquer approach of [37] does not give the desired running time for us (it loses another logarithmic factor in the exponent). To get around this problem, we extend Marx and Pilipczuk’s technique of noose-based separators to a noose-based dynamic programming algorithm. Additionally, we introduce well-spaced and valid sphere cut decompositions, and prove that they are in one-to-one correspondence with independent sets of $G$.
Finally, we consider approximation algorithms for Maximum Independent Set when the graph in question has low ply, that is, each point of $\mathbb{H}^{2}$ is covered by at most $\ell$ disks. We show the following.
Theorem 1.6.
@@
An important component of the algorithm’s correctness proof is bounding the so-called degeneracy of hyperbolic uniform disk graphs in terms of their clique size and in terms of their ply. This generalizes the Euclidean results of [34]. When compared to the $n^{\mathcal{O}(1/\varepsilon)}$ algorithm for disk graphs by Chan [13] or the $2^{\mathcal{O}(1/\varepsilon)}n$ algorithm in planar graphs by Baker [2], both of which are conditionally optimal [35], our algorithm has a surprising quasi-polynomial dependence on $1/\varepsilon$ rather than exponential. The dependence on $r$ exhibits the classic Euclidean-to-hyperbolic scaling seen in our earlier structural results. For $\varepsilon=1/n$, we are guaranteed to get the optimal solution, and setting $\ell=n$ covers the general case. The resulting running time for $1/\varepsilon=\ell=n$ is $n^{\mathcal{O}\left(1+\frac{1}{r}\log n\right)}$, which matches our exact algorithm. Consequently, our approximation scheme is a generalization of our exact algorithm.
Outline.
The remainder of this article is structured as follows. After the introduction of important definitions and notation in Section 2, we first give a high-level overview of our results in Section 3. The subsequent sections then discuss all of our results in more detail. Finally, we conclude with a short summary and discussion of future directions in Section 10.
@@ 2 Preliminaries
An independent set in a graph is a set of pairwise non-adjacent vertices. In the Independent Set problem, we are given [^7] In this article, the graphs are always intersection graphs given through their geometric representation. a graph $G$ and a number $k$, and the goal is to decide if there is an independent set of size $k$. In Maximum Independent Set, the input is a graph $G$, and the goal is to output the maximum value $k$ such that $G$ has an independent set of size $k$.
Let $G=(V,E)$ be a graph. A vertex set $S\subseteq V$ is a separator if $V$ can be partitioned into non-empty subsets $S$, $V_{1}$, and $V_{2}$ such that there is no edge between $V_{1}$ and $V_{2}$, i.e., for every $v_{1}\in V_{1}$ and $v_{2}\in V_{2}$, it holds that $\{v_{1},v_{2}\}\notin E$. The separator is $\beta$-balanced if $|V_{1}|,|V_{2}|\leq\beta n$ for some $\beta<1$.
@@
We will need the notion of weighted treewidth [45]. Here each vertex has a weight, and the weighted width of a tree decomposition is the maximum over the bags of the sum of the weights of the vertices in the bag (note: without the minus $1$). The weighted treewidth of a graph is the minimum weighted width over its tree decompositions. Now let $\mathcal{P}$ be a clique partition of a given graph $G$. We apply the concept of weighted treewidth to $G_{\mathcal{P}}$, where we assign each vertex $C$ of $G_{\mathcal{P}}$ the weight $\log(|C|+1)$, and refer to this weighting whenever we talk about the weighted treewidth of a contraction $G_{\mathcal{P}}$. For any given $\mathcal{P}$, the weighted treewidth of $G_{\mathcal{P}}$ with the above weighting is referred to as the $\mathcal{P}$-flattened treewidth of $G$.
@@
A polygon is a cycle together with a planar drawing in which all edges are drawn as line segments. Note that this enables the use of graph notation, e.g., using $V(P)$ to refer to the vertices of polygon $P$.
@@
Figure 3: Part (a) shows a visualization of the Poincaré disk including (parallel) lines, a circle with its center, a triangle and generalized polygons. Part (b) shows a visualization of the Beltrami-Klein model including (parallel) lines, a triangle, and a generalized polygon. Part (c) shows a set of sites $S\subseteq\mathbb{H}^{2}$ in the Poincaré disk model, $\mathcal{D}(S)$ (black), and $\mathcal{V}(S)$ (gray); Voronoi vertices are filled squares, ideal Voronoi vertices are empty squares.
@@
We write $\mathbb{H}^{2}$ to refer to the hyperbolic plane and $\mathbb{R}^{2}$ to refer to the Euclidean plane. The Poincaré disk is a model of $\mathbb{H}^{2}$ that maps the whole hyperbolic plane into the interior of a Euclidean unit disk; see Figure 3. We refer to the center of the Poincaré disk as origin. Straight lines in $\mathbb{H}^{2}$ are represented as circular arcs perpendicular to the boundary of the Poincaré disk or as chords through the origin. Hyperbolic circles are represented as Euclidean circles. The center of the hyperbolic circle is, however, farther from the origin than the Euclidean center of its representation. The Poincaré disk is conformal, i.e., angle preserving.
@@
Points on the boundary of the Poincaré disk are called ideal points. These are not part of the hyperbolic plane, but form a boundary of infinitely far points. An ideal arc between two ideal points $q_{1}$ and $q_{2}$ is the set of ideal points between $q_{1}$ and $q_{2}$ when moving clockwise on the boundary of the Poincaré disk. A generalized polygon is a cycle together with a planar drawing that maps each vertex to a point in $\mathbb{H}^{2}$ or to an ideal point. Each edge is either a line segments between points, a ray from a point to an ideal point, a line between two ideal points, or an ideal arc between two ideal points. Note that the vertices do not completely determine the generalized polygon, as two ideal points can be connected via a line or an ideal arc. Moreover, a 2-cycle can yield a generalized polygon by mapping both vertices to ideal points and one edge to the line and the other to the ideal arc between the two. Figure 3 shows two generalized polygons with their interior shaded in blue.
@@
For a straight line $\ell$ and a point $p\notin\ell$, there are infinitely many parallel lines through $p$, i.e., lines through $p$ that do not intersect $\ell$. Two of these parallel lines are special in the sense that they are the closest to not being parallel. They each share an ideal endpoint with $\ell$ and are called limiting parallels. Let $q\in\ell$ be such that $pq$ is perpendicular to $\ell$. Then the angle between $pq$ and the two limiting parallels is the same on both sides and only depends on the length $x=|pq|$. It is called the angle of parallelism and denoted by $\Pi(x)$. It holds that $\sin(\Pi(x))=1/\cosh(x)$ [20, page 402].
@@
Let $a$ be the sum of interior angles of a hyperbolic triangle. Then $a<\pi$ and the area of the triangle is $\pi-a$. Note that this implies that the area of a triangle is upper bounded by $\pi$. More generally, the area of a $k$-gon is bounded by $(k-2)\pi$. The area of a disk with hyperbolic radius $r$ is $4\pi\sinh^{2}(r/2)$, which is in $\Theta(e^{r})$ for $r\to\infty$ and in $\Theta(r^{2})$ for $r\to 0$.
@@
The Beltrami–Klein model also uses the interior of the Euclidean unit disk as ground space. Hyperbolic straight lines are represented as chords of the unit disk, see Figure 3. This model is not conformal, but enables an easy translation of Euclidean line arrangements into the hyperbolic plane.
@@
Let $S$ be a set of sites, i.e., a set of designated points in $\mathbb{H}^{2}$. The Voronoi cell of a site $s\in S$ is the set of points that are closer to $s$ than to any other site. When considered in the Poincaré disk, the boundary of each Voronoi cell is a generalized polygon [^8] We note that Voronoi cells containing ideal points in their boundary are actually unbounded in the hyperbolic plane.. The non-ideal segments are part of the perpendicular bisector of two sites and are called Voronoi edges; they are illustrated in blue in Figure 3. The points in which three or more Voronoi edges meet are called Voronoi vertices. The ideal points in which unbounded Voronoi edges end are called ideal Voronoi vertices. We define the Voronoi diagram $\mathcal{V}(S)$ of $S$ as the following plane graph. Its vertex set is the set of all (ideal) Voronoi vertices. The edge set of $\mathcal{V}(S)$ is comprised of the Voronoi edges and the set of ideal arcs connecting consecutive ideal vertices (red in Figure 3).
The weak dual of the Voronoi diagram is called Delaunay complex and denoted by $\mathcal{D}(S)$; it is illustrated in black in Figure 3. The edges of $\mathcal{D}(S)$ are exactly the edges dual to the Voronoi edges, i.e., the edges of $\mathcal{V}(S)$ that are not ideal arcs. Thus, sites $s_{1}\in S$ and $s_{2}\in S$ are connected in the $\mathcal{D}(S)$ if and only if their Voronoi cells share a boundary. In contrast to the Euclidean case, the outer face of $\mathcal{D}(S)$ is not the convex hull of $S$. In fact, its boundary is not necessarily a simple cycle; see Figure 3. A site is an outer vertex of $\mathcal{D}(S)$ if and only if the corresponding Voronoi cell is unbounded.